40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6252     Entry time: Fri Feb 3 19:57:26 2012
Author: rana 
Type: Configuration 
Category: SUS 
Subject: PSL Shutter closed for SUS hysteresis test 

Fri Feb 03 19:57:20 2012

Fri Feb 03 20:25:19 2012 :   Aligned all SUS to center their OL beams

Fri Feb 03 20:29:21 2012:    Aligned all SUS to make OL_PIT = 0.5

ELOG V3.1.3-